OpenStudy (anonymous):

Simplify. (-1/2)(4x)(-2)(7y)(-1) A. -28xy B. -28 C. 28xy D. 27 Plz help!

OpenStudy (jhannybean):

\[-\frac{1}{2} \cdot 4x \cdot -2 \cdot 7y \cdot -1\]First solve \(-\dfrac{1}{2} \cdot 4x\)

OpenStudy (jackthegreatest):

-28xy

OpenStudy (anonymous):

Thanks to both of you:)

OpenStudy (jackthegreatest):

combine the like terms